OpenStudy (anonymous):

In a circle of radius 25 inches, a central angle of 35 will intersect the circle forming an arc of length ____. PLEASE HELP

Parth (parthkohli):

\[\rm arc ~length = circumference \times \frac{\theta}{360}\]

OpenStudy (anonymous):

can you show me how to get the answer?

Parth (parthkohli):

Hmm -- the circumference is \(2\pi r = 2 \times \pi \times 5 = 10\pi\). \(\theta\) means the central angle, i.e., \(35^{\circ}\)

Parth (parthkohli):

That formula will make much more sense to you with a diagram.

Parth (parthkohli):



OpenStudy (anonymous):

ok but whats the answer?

OpenStudy (mathmale):

I'd prefer to start out with the simplest formula relating arc length, s, to radius, r, and central angle, theta: s = r theta, where theta MUST be in radians.